An object has a weight of 690 Newton on earth, specify (a) its mass in kilograms, (b) its weight in pound-force, and (c) its mass in slugs. If the object is on other celestial objects, where the gravitational acceleration is g = 5.50 ft/s2 , determine (d) its mass in kilograms, and (e) its weight in pounds-force.


1
Expert's answer
2021-04-01T18:42:30-0400

a)


"m=\frac{690}{9.8}=70.4\ kg"

b)


"W=\frac{690}{4.448}=155\ lb_F"

c)


"m=\frac{155}{32.17405}=4.82\ slugs"

d)


"m=70.4\ kg"

e)


"W'=4.82(5.5)=26.5\ lb_f"